### v1

# An uncountable subring of $\mathbb{R}$ with Hausdorff dimension zero

Abstract.
We construct an uncountable subring of $\mathbb{R}$ with Hausdorff dimension zero (and hence of Lebesgue measure zero).
We denote by $\mathbb{N}_{0}$ the set of non-negative integers, by $\mathbb{N}$ the set of positive integers and by $\mathbb{Z}$ the set of integers. Let
$S:=\{0\}\cup\{2^{n}:n\in\mathbb{N}_{0}\}=\{0,1,2,2^{2},2^{3},\ldots\}.$
For every $n\in\mathbb{N}$ by $nS$, we denote the $n^{\text{th}}$ sumset of $S$, i.e.,
$nS:=\{s_{1}+\cdots+s_{n}:s_{1},\ldots,s_{n}\in S\}.$
For every $n\in\mathbb{N}$ we define
$A_{n}:=\left\{\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}:(x_{k})_{k\in nS}\text{is a bounded sequence of integers}\right\}.$
Theorem 1.
Let $n\in\mathbb{N}$. The set $A_{n}$ is an uncountable subgroup of $\mathbb{R}$.
Proof.
We first show that $A_{n}$ is a subgroup of $\mathbb{R}$. Trivially, $0\in A_{n}$ and $-x\in A_{n}$ whenever $x\in A_{n}$. Thus, it suffices to show that $A_{n}$ is closed under addition. Let $x,y\in A_{n}$. Therefore,
$x=\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}\quad\mbox{and}\quad y=\sum\limits_{k\in nS}\frac{y_{k}}{2^{k}}$
where $(x_{k})_{k\in nS}$ and $(y_{k})_{k\in nS}$ are bounded sequences of integers. It follows that
$x+y=\sum\limits_{k\in nS}\frac{x_{k}+y_{k}}{2^{k}}$
is also an element of $A_{n}$.
Clearly, the set $A_{n}$ is uncountable, because by the uniqueness of binary expansions which do not end with an infinite string of $1$’s, the set of real numbers of the form
$\sum\limits_{k\in nS}\frac{z_{k}}{2^{k}}\quad\mbox{with }z_{k}\in\{0,1\}$
is an uncountable subset of $A_{n}$. ∎
We now proceed to show that the set $A_{n}$ has Hausdorff dimension zero for every $n\in\mathbb{N}$. For every $n\in\mathbb{N}$ write
$A_{n}:=\bigcup_{t=1}^{\infty}A_{n,t}$
where for every $t\in\mathbb{N}$ we let
$A_{n,t}:=\left\{\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}\in\mathbb{R}:x_{k}\in\mathbb{Z}\mbox{ and }|x_{k}|\leq t\mbox{ for all }k\in nS\right\}.$
We will need the following result to show that the set $A_{n,t}$ has Hausdorff dimension zero for every $n,t\in\mathbb{N}$.
Lemma 1.
For $n\in\mathbb{N}$ and $z\geq 1$, define
$g_{n}(z):=\emph{{\#}}\{k\in nS:k\leq z\}.$
Then we have the bound
$g_{n}(z)\leq(2+\log_{2}z)^{n}.$
Proof.
Clearly, for every $z\geq 1$, we have
$g_{n}(z)\leq g_{1}(z)^{n}\mbox{ and }g_{1}(z)\leq 2+\text{log}_{2}z.$
The result follows. ∎
Theorem 2.
The set $A_{n,t}$ has Hausdorff dimension zero for all $n,t\in\mathbb{N}$.
Proof.
Write $nS=\{k_{m}:m\in\mathbb{N}\}$ where $(k_{m})_{m\in\mathbb{N}}$ is an increasing sequence. Let $x\in nS$ and write
$x=\sum\limits_{i=1}^{\infty}\frac{a_{i}}{2^{k_{i}}}$
with $a_{i}\in\mathbb{Z}$ and $|a_{i}|\leq t$. Pick $j\in\mathbb{N}$ and split $x$ into
$x=m_{x}+e_{x},$
where
$m_{x}:=\sum\limits_{i=1}^{j-1}\frac{a_{i}}{2^{k_{i}}}\quad\mbox{and}\quad e_{x}:=\sum\limits_{i=j}^{\infty}\frac{a_{i}}{2^{k_{i}}}.$
Then using the triangle inequality, we have the bound
$|e_{x}|\leq t\sum\limits_{i=j}^{\infty}\frac{1}{2^{k_{i}}}\leq t\sum\limits_{h=k_{j}}^{\infty}\frac{1}{2^{h}}=\frac{t}{2^{k_{j}-1}}.$
Let $r_{j}=t/2^{k_{j}-1}$. It follows that
$A_{n,t}\subseteq\bigcup_{m\in X}I_{m},$
where
$X:=\left\{\sum\limits_{i=1}^{j-1}\frac{a_{i}}{2^{k_{i}}}:a_{i}\in\mathbb{Z},\ |a_{i}|\leq t\right\}$
and
$I_{m}:=[m-r_{j},m+r_{j}].$
Then #$\emph{{\#}}(X)\leq(2t+1)^{j-1}$ and for every $m\in X$ we see that $\mu(I_{m})=2r_{j}$ where $\mu(I_{m})$ is the length of the interval $I_{m}$.
Let $d>0$ be a real number, let $\delta_{j}:=4r_{j}$, and let $\mathcal{H}_{\delta_{j}}^{d}(A_{n,t})$ be defined as in [2]. We see that
$\mathcal{H}_{\delta_{j}}^{d}(A_{n,t})\leq\sum\limits_{m\in X}\mu(I_{m})^{d}=\emph{{\#}}(X)(2r_{j})^{d}\leq\frac{(2t+1)^{j-1}t^{d}}{2^{(k_{j}-2)d}}.$
We claim that
(1) $\lim\limits_{j\rightarrow\infty}\frac{(2t+1)^{j-1}t^{d}}{2^{(k_{j}-2)d}}=0.$
Since $\delta_{j}\to 0$ as $j\to\infty$, this implies that $\mathcal{H}^{d}(A_{n,t})=0$ where $\mathcal{H}^{d}(A_{n,t})$ is the $d$-dimensional Hausdorff measure of $A_{n,t}$. (For the definition of this measure and related terms, see [2].) Since this holds for all $d>0$, we conclude that the Hausdorff dimension of $A_{n,t}$ is zero.
Equation (1) above is seen as follows. First, it is convenient to note that (1) holds if
$\lim\limits_{j\rightarrow\infty}\frac{(3t)^{j}}{2^{k_{j}d}}=0.$
By taking logarithms, we see that (1) holds if
$\lim\limits_{j\rightarrow\infty}\big{(}k_{j}d-j\log_{2}(3t)\big{)}=\infty.$
As $d>0$, we let $c$ be the constant $(1/d)\log_{2}(3t)$ and we see that (1) holds if
(2) $\lim\limits_{j\to\infty}(k_{j}-cj)=\infty.$
Now using Lemma 1, we have
$j=g_{n}(k_{j})\leq(2+\log_{2}k_{j})^{n},$
which implies
$k_{j}\geq 2^{\sqrt[n]{j}-2}.$
Hence, (2) follows, which justifies our claim. ∎
Let $n\in\mathbb{N}$. From Theorems 1 and 2 it follows that the set $A_{n}$ is an uncountable subgroup of $\mathbb{R}$ of Hausdorff dimension zero. As $0\in S$, we have $nS\subseteq(n+1)S$, and hence $A_{n}\subseteq A_{n+1}$. Let
$A:=\bigcup_{n=1}^{\infty}A_{n}.$
We deduce that the set $A$ is an uncountable subgroup of $\mathbb{R}$ of Hausdorff dimension zero. We aim to establish that $A$ is in fact a subring. To this end, we need the following result.
Lemma 2.
Let $n\in\mathbb{N}$ and let $a_{n}(s)$ be the number of representations of an element $s\in\mathbb{N}$ as a sum $s=t_{1}+t_{2}+\cdots+t_{n}$ of $n$ elements $t_{1},t_{2},\ldots,t_{n}\in S$. Then there exists $c_{n}\in\mathbb{R}$ such that $a_{n}(s)\leq c_{n}$ for every $s\in\mathbb{N}$.
Proof.
Clearly $a_{1}(s)\leq 1$ for every $s\in\mathbb{N}$ and hence, we may take $c_{1}:=1$. Now assume that $n\geq 2$ and that there exists $c_{n-1}\in\mathbb{R}$ such that $a_{n-1}(s)\leq c_{n-1}$ for every $s\in\mathbb{N}$. Let $s\in\mathbb{N}$ and suppose
$s=t_{1}+\cdots+t_{n}$
is a representation of $s$ as a sum of $n$ elements of $S$. Then for at least one $i\in\{1,\ldots,n\}$, we have $t_{i}\geq s/n$ and hence $t_{i}\neq 0$. So there exists $k\in\mathbb{N}_{0}$ such that $t_{i}=2^{k}$. As $s/n\leq t_{i}\leq s$, we observe that $\log_{2}s-\log_{2}n\leq k\leq\log_{2}s$. So if
$B_{n,s}:=[\,\log_{2}s-\log_{2}n,\,\log_{2}s\,]\cap\mathbb{N}_{0},$
then we see that
$a_{n}(s)\leq n\sum\limits_{k\in B_{n,s}}a_{n-1}(s-2^{k})\leq n(1+\log_{2}n)c_{n-1}.$
So we are done by taking $c_{n}:=n(1+\log_{2}n)c_{n-1}$. ∎
Theorem 3.
The set $A$ is a a subring of $\mathbb{R}$.
Proof.
It suffices to show that $1\in A$ and that $A$ is closed under multiplication. We see that
$1=\frac{1}{2^{0}}+\sum\limits_{k\in S\setminus\{0\}}\frac{0}{2^{k}}\in A_{1}.$
To establish that $A$ is closed under multiplication, we proceed as follows. Let $x,y\in A$. There exists $m,n\in\mathbb{N}$ such that $x\in A_{m}$ and $y\in A_{n}$. Therefore,
$x=\sum\limits_{k\in mS}\frac{x_{k}}{2^{k}}\quad\mbox{and}\quad y=\sum\limits_{\ell\in nS}\frac{y_{\ell}}{2^{\ell}}$
where $(x_{k})_{k\in mS}$ and $(y_{\ell})_{\ell\in nS}$ are bounded sequences of integers. Now
$xy=\left(\sum\limits_{k\in mS}\frac{x_{k}}{2^{k}}\right)\left(\sum\limits_{\ell\in nS}\frac{y_{\ell}}{2^{\ell}}\right)=\sum\limits_{(k,\ell)\in mS\times nS}\frac{x_{k}y_{\ell}}{2^{k+\ell}}=\sum\limits_{r\in(m+n)S}\frac{z_{r}}{2^{r}}$
where
$z_{r}:=\sum\limits_{\begin{subarray}{c}(k,\ell)\in mS\times nS\\ k+\ell=r\end{subarray}}x_{k}y_{\ell}.$
The number of representations of an element $r\in(m+n)S$ as a sum $r=k+\ell$ of elements $k\in mS$ and $\ell\in nS$ is bounded by the number of representations of $r$ which are of the form $r=t_{1}+t_{2}+\cdots+t_{m+n}$ where $t_{1},t_{2},\ldots,t_{m+n}\in S$. Hence, we deduce that the sequence $(z_{r})_{r\in(m+n)S}$ is bounded as a consequence of Lemma 2 and the boundedness of the sequences $(x_{k})_{k\in mS}$ and $(y_{\ell})_{\ell\in nS}$. It follows that $xy\in A_{m+n}$. This establishes that $A$ is closed under multiplication. ∎
We now proceed to show that the only rational numbers in $A$ are the dyadic rationals. In particular, it will follow that the ring $A$ is not a field.
Lemma 3.
Let $r,n\in\mathbb{N}$. Write $nS=\{k_{m}:m\in\mathbb{N}\}$ where $(k_{m})_{m\in\mathbb{N}}$ is an increasing sequence. Then there exists $m\in\mathbb{N}$ such that $k_{m+1}-k_{m}>r$.
Proof.
For every $z\in\mathbb{N}$, let $g_{n}(z):=\emph{{\#}}\{k\in nS:k\leq z\}$. If the conclusion is false, then $g_{n}(z)\geq z/r$ for every $z\in\mathbb{N}$. This contradicts Lemma 1. ∎
Proposition 1.
Let $x\in A$ and $l\in\mathbb{N}$. Then there exists a binary expansion of $x$ which contains either a string of zeroes or a string of ones having length $l$.
Proof.
Let $x\in A_{n,t}$. Then there exists a sequence of integers $(x_{k})_{k\in nS}$ with $|x_{k}|\leq t$ for all $k\in nS$ such that
$x=\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}=\sum\limits_{i=1}^{\infty}\frac{x_{k_{i}}}{2^{k_{i}}}.$
Let $m\in\mathbb{N}$. We observe that
$2^{k_{m}}x=q_{m}+r_{m}$
where
$q_{m}:=\sum\limits_{i=0}^{m}2^{k_{m}-k_{i}}x_{k_{i}}\mbox{ and }r_{m}:=\sum\limits_{i=m+1}^{\infty}\frac{x_{k_{i}}}{2^{k_{i}-k_{m}}}.$
Then $q_{m}\in\mathbb{Z}$ and we have
$\left|r_{m}\right|\leq~{}t\sum\limits_{j=k_{m+1}-k_{m}}^{\infty}\frac{1}{2^{j}}=\frac{2t}{2^{k_{m+1}-k_{m}}}.$
By Lemma 3, there exists $m\in\mathbb{N}$ such that $2^{k_{m+1}-k_{m}}>2^{l+1}t$ and so $|r_{m}|<1/2^{l}$. Hence,
$\text{frac}\left(2^{k_{m}}x\right)=\begin{cases}r_{m},&\mbox{ if }r_{m}\in\left[\,0,\frac{1}{2^{l}}\right);\\ 1+r_{m},&\mbox{ if }r_{m}\in\left(-\frac{1}{2^{l}},0\right);\end{cases}$
and so
$\text{frac}(2^{k_{m}}x)\in\left[0,\tfrac{1}{2^{l}}\right)\cup\left(1-\tfrac{1}{2^{l}},1\right),$
where frac$(z):=z-[z]$ is the fractional part of $z\in\mathbb{R}$. It follows that the digits after the dot in a binary expansion of $x$ from the $(k_{m}+1)^{\text{th}}$ position to the $(k_{m}+l)^{\text{th}}$ position are all equal. ∎
Corollary 1.
A number $x\in A$ is rational if and only if there exists $a\in\mathbb{Z}$ and $k\in\mathbb{N}_{0}$ such that $x=a/2^{k}$.
Proof.
From Proposition 1, we see that the ring $A$ only contains those rational numbers which have a finite binary expansion. The result follows. ∎
Remark 1.
The above construction of an uncountable subring of $\mathbb{R}$ of Hausdorff dimension zero goes through if the initial set $S=\{0,1,2,2^{2},2^{3},\ldots\}$ is replaced by a set $T\cup\{0\}$ where $T$ is a subset of $\mathbb{N}$ which has the following property: There exists $b\in\mathbb{R}$ such that for every $n\in\mathbb{N}$, the cardinality of the set $\{\log t:t\in T\}\cap[n,n+1)$ is less than $b$.
Acknowledgment: We thank Prof. Shrikrishna G. Dani for sharing with us a preprint of [1] in which we see a different construction of an uncountable subgroup of $\mathbb{R}$ having Hausdorff dimension zero.

### v2

@@
 Abstract.
-We construct an uncountable subring of $\mathbb{R}$ with Hausdorff dimension zero (and hence of Lebesgue measure zero).
-We denote by $\mathbb{N}_{0}$ the set of non-negative integers, by $\mathbb{N}$ the set of positive integers and by $\mathbb{Z}$ the set of integers. Let
-$S:=\{0\}\cup\{2^{n}:n\in\mathbb{N}_{0}\}=\{0,1,2,2^{2},2^{3},\ldots\}.$
-For every $n\in\mathbb{N}$ by $nS$, we denote the $n^{\text{th}}$ sumset of $S$, i.e.,
+We construct a subring as mentioned in the title (hence this subring has Lebesgue measure zero).
+1. Preliminaries
+For a subset $B\subseteq\mathbb{R}$ we let $d(B)$ denote the diameter of $B$. Given a countable collection of subsets $\mathcal{C}\subseteq\mathcal{P}(\mathbb{R})$ and $s\geq 0$ we define
+$d(s,\mathcal{C}):=\sum_{B\in\mathcal{C}}d(B)^{s}.$
+We note that $d(s,\mathcal{C})\in[0,\infty]$ and does not depend on the ordering of the subsets in the collection $\mathcal{C}$.
+For $\delta>0$, we say that a collection of subsets $\mathcal{C}$ is a $\delta$-cover of a subset $F\subseteq\mathbb{R}$ if for every $B\in\mathcal{C}$ we have $d(B)\leq\delta$ and if $F\subseteq\cup\,\mathcal{C}$.
+Given a subset $F\subseteq\mathbb{R}$ and $s,\delta\in\mathbb{R}$ with $s\geq 0$, and $\delta>0$, we define
+$\mathcal{H}^{s}_{\delta}(F)=\inf\big{\{}d(s,\mathcal{C}):\mathcal{C}\mbox{ is a countable $\delta$-cover of $F$}\big{\}}.$
+If $\delta_{1}<\delta_{2}$, then we see that $\mathcal{H}^{s}_{\delta_{1}}(F)\geq\mathcal{H}^{s}_{\delta_{2}}(F)$.
+The $s$-dimensional Hausdorff measure $\mathcal{H}^{s}(F)$ is defined to be the limit of $\mathcal{H}_{\delta}^{s}(F)$ as $\delta\to 0^{+}$.
+It can be shown (see [4]) that there exists $\alpha\in[0,1]$ such that for every $s\in[0,\alpha)$ we have $\mathcal{H}^{s}(F)=\infty$, and for every $s>\alpha$ we have $\mathcal{H}^{s}(F)=0$.
+The number $\alpha$ is called the Hausdorff dimension of $F$ and is denoted by $\dim_{\mathcal{H}}F$.
+If $F=\cup\,\mathcal{C}$ where $\mathcal{C}$ is a countable collection of subsets of $\mathbb{R}$, then $\dim_{\mathcal{H}}F=\sup\big{\{}\dim_{\mathcal{H}}A:A\in\mathcal{C}\big{\}}$.
+2. Construction
+We give an explicit construction of a proper subring of $\mathbb{R}$ which is a Borel set. We also show explicitly that it has Hausdorff dimension zero.
+We let $\mathbb{N}_{0}$ denote the set of non-negative integers, $\mathbb{N}$ denote the set of positive integers, and $\mathbb{Z}$ denote the set of integers.
+For every $a,b\in\mathbb{Z}$ we let $a,b]_{\mathbb{Z}}=\{k\in\mathbb{Z}:a\leq k\leq b\}$.
+Let
+$S:=\{0,1,2,2^{2},2^{3},\ldots\}=\{0\}\cup\{2^{n}:n\in\mathbb{N}_{0}\}.$
+For every $n\in\mathbb{N}$ we let the $n^{\text{th}}$ sumset of $S$ be defined as
 $nS:=\{s_{1}+\cdots+s_{n}:s_{1},\ldots,s_{n}\in S\}.$
 For every $n\in\mathbb{N}$ we define
-$A_{n}:=\left\{\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}:(x_{k})_{k\in nS}\text{is a bounded sequence of integers}\right\}.$
+[…]
 Theorem 1.
@@
 We first show that $A_{n}$ is a subgroup of $\mathbb{R}$.
-Trivially, $0\in A_{n}$ and $-x\in A_{n}$ whenever $x\in A_{n}$.
+Trivially, we have $0\in A_{n}$, and that $x\in A_{n}$ whenever $x\in A_{n}$.
 Thus, it suffices to show that $A_{n}$ is closed under addition.
 Let $x,y\in A_{n}$.
-Therefore,
-$x=\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}\quad\mbox{and}\quad y=\sum\limits_{k\in nS}\frac{y_{k}}{2^{k}}$
-where $(x_{k})_{k\in nS}$ and $(y_{k})_{k\in nS}$ are bounded sequences of integers. It follows that
+Therefore, there exists $t\in\mathbb{N}$ and for every $k\in nS$ there exist $x_{k},y_{k}\in[-t,t\,]_{\mathbb{Z}}$ such that
+$x=\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}\quad\mbox{and}\quad y=\sum\limits_{k\in nS}\frac{y_{k}}{2^{k}}.$
+It follows that
 $x+y=\sum\limits_{k\in nS}\frac{x_{k}+y_{k}}{2^{k}}$
-is also an element of $A_{n}$.
-Clearly, the set $A_{n}$ is uncountable, because by the uniqueness of binary expansions which do not end with an infinite string of $1$’s, the set of real numbers of the form
-$\sum\limits_{k\in nS}\frac{z_{k}}{2^{k}}\quad\mbox{with }z_{k}\in\{0,1\}$
+is also an element of $A_{n}$ since $x_{k}+y_{k}\in[-2t,2t\,]_{\mathbb{Z}}$ for every $k\in nS$.
+Also, the set $A_{n}$ is uncountable, because by the uniqueness of binary expansions which do not end with an infinite string of ones, the set of real numbers of the form
+[…]
 is an uncountable subset of $A_{n}$. ∎
@@
 where for every $t\in\mathbb{N}$ we let
-$A_{n,t}:=\left\{\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}\in\mathbb{R}:x_{k}\in\mathbb{Z}\mbox{ and }|x_{k}|\leq t\mbox{ for all }k\in nS\right\}.$
+$A_{n,t}:=\left\{\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}\in\mathbb{R}:\mbox{ for every }k\in nS\mbox{ we have }x_{k}\in[-t,t\,]_{\mathbb{Z}}\right\}.$
+For every $n,\ell\in\mathbb{N}$ we let $(nS)_{\leq\ell}=\{k\in nS:k\leq\ell\}$ and we let $(nS)_{>\ell}=\{k\in nS:k>\ell\}$.
 We will need the following result to show that the set $A_{n,t}$ has Hausdorff dimension zero for every $n,t\in\mathbb{N}$.
 Lemma 1.
-For $n\in\mathbb{N}$ and $z\geq 1$, define
-$g_{n}(z):=\emph{{\#}}\{k\in nS:k\leq z\}.$
+For every $n,\ell\in\mathbb{N}$ we define
+$g_{n}(\ell):=\emph{{\#}}(nS)_{\leq\ell}.$
 Then we have the bound
-$g_{n}(z)\leq(2+\log_{2}z)^{n}.$
+$g_{n}(\ell)\leq(2+\log_{2}\ell)^{n}.$
 Proof.
-Clearly, for every $z\geq 1$, we have
-$g_{n}(z)\leq g_{1}(z)^{n}\mbox{ and }g_{1}(z)\leq 2+\text{log}_{2}z.$
+We observe that for every $z\in\mathbb{N}$, we have
+$g_{n}(\ell)\leq g_{1}(\ell)^{n}\mbox{ and }g_{1}(\ell)\leq 2+\log_{2}\ell.$
 The result follows. ∎
@@
 Proof.
-Write $nS=\{k_{m}:m\in\mathbb{N}\}$ where $(k_{m})_{m\in\mathbb{N}}$ is an increasing sequence. Let $x\in nS$ and write
-$x=\sum\limits_{i=1}^{\infty}\frac{a_{i}}{2^{k_{i}}}$
-with $a_{i}\in\mathbb{Z}$ and $|a_{i}|\leq t$. Pick $j\in\mathbb{N}$ and split $x$ into
-$x=m_{x}+e_{x},$
+Let $x\in A_{n,t}$. Then for every $k\in nS$ there exist $a_{k}\in[-t,t\,]_{\mathbb{Z}}$ such that
+$x=\sum\limits_{k\in nS}\frac{a_{k}}{2^{k}}.$
+Pick $\ell\in\mathbb{N}$ and write $x$ as
+$x=y+z,$
 where
-$m_{x}:=\sum\limits_{i=1}^{j-1}\frac{a_{i}}{2^{k_{i}}}\quad\mbox{and}\quad e_{x}:=\sum\limits_{i=j}^{\infty}\frac{a_{i}}{2^{k_{i}}}.$
+$y:=\sum\limits_{k\,\in\,(nS)_{\leq\ell}}\frac{a_{k}}{2^{k}}\quad\mbox{and}\quad z:=\sum\limits_{k\,\in\,(nS)_{>\ell}}\frac{a_{k}}{2^{k}}.$
 Then using the triangle inequality, we have the bound
-$|e_{x}|\leq t\sum\limits_{i=j}^{\infty}\frac{1}{2^{k_{i}}}\leq t\sum\limits_{h=k_{j}}^{\infty}\frac{1}{2^{h}}=\frac{t}{2^{k_{j}-1}}.$
-Let $r_{j}=t/2^{k_{j}-1}$. It follows that
-$A_{n,t}\subseteq\bigcup_{m\in X}I_{m},$
-where
-$X:=\left\{\sum\limits_{i=1}^{j-1}\frac{a_{i}}{2^{k_{i}}}:a_{i}\in\mathbb{Z},\ |a_{i}|\leq t\right\}$
-and
-$I_{m}:=[m-r_{j},m+r_{j}].$
-Then #$\emph{{\#}}(X)\leq(2t+1)^{j-1}$ and for every $m\in X$ we see that $\mu(I_{m})=2r_{j}$ where $\mu(I_{m})$ is the length of the interval $I_{m}$.
-Let $d>0$ be a real number, let $\delta_{j}:=4r_{j}$, and let $\mathcal{H}_{\delta_{j}}^{d}(A_{n,t})$ be defined as in [2]. We see that
-$\mathcal{H}_{\delta_{j}}^{d}(A_{n,t})\leq\sum\limits_{m\in X}\mu(I_{m})^{d}=\emph{{\#}}(X)(2r_{j})^{d}\leq\frac{(2t+1)^{j-1}t^{d}}{2^{(k_{j}-2)d}}.$
+$|z|\leq t\sum\limits_{k\,\in\,(nS)_{>\ell}}\frac{1}{2^{k}}\leq t\sum\limits_{k=\ell+1}^{\infty}\frac{1}{2^{k}}=\frac{t}{2^{\ell}}.$
+Let
+$B:=\left\{\sum\limits_{k\,\in\,(nS)_{\leq\ell}}\frac{a_{k}}{2^{k}}:\mbox{ for every }k\in(nS)_{\leq\ell}\mbox{ we have }a_{k}\in[-t,t\,]_{\mathbb{Z}}\right\}$
+and let $r=t/2^{\ell}$. It follows that
+$A_{n,t}\subseteq\bigcup_{b\in B}I_{b}\mbox{ where }I_{b}=[b-r,b+r]\mbox{ for every }b\in B.$
+Let $c_{\ell}=\emph{{\#}}\,(nS)_{\leq\ell}$.
+Then $\emph{{\#}}B\leq(2t+1)^{c_{\ell}}$ and for every $b\in B$ we see that $\mu(I_{b})=2r$ where $\mu$ is the Lebesgue measure on $\mathbb{R}$.
+Let $s\in(0,1]$.
+We see that
+$\mathcal{H}_{2r_{\ell}}^{s}(A_{n,t})\leq\sum\limits_{b\in B}\mu(I_{b})^{s}=\emph{{\#}}B\,(2r)^{s}\leq(2t+1)^{c_{\ell}}\frac{t^{s}}{2^{(\ell-1)s}}.$
 We claim that
-(1) $\lim\limits_{j\rightarrow\infty}\frac{(2t+1)^{j-1}t^{d}}{2^{(k_{j}-2)d}}=0.$
-Since $\delta_{j}\to 0$ as $j\to\infty$, this implies that $\mathcal{H}^{d}(A_{n,t})=0$ where $\mathcal{H}^{d}(A_{n,t})$ is the $d$-dimensional Hausdorff measure of $A_{n,t}$.
-(For the definition of this measure and related terms, see [2].) Since this holds for all $d>0$, we conclude that the Hausdorff dimension of $A_{n,t}$ is zero.
+(1) $\lim\limits_{\ell\rightarrow\infty}(2t+1)^{c_{\ell}}\frac{t^{s}}{2^{(\ell-1)s}}=0.$
+Since $r\to 0$ as $\ell\to\infty$, this implies that $\mathcal{H}^{s}(A_{n,t})=0$. As this holds for all $s\in(0,1]$, we conclude that the Hausdorff dimension of $A_{n,t}$ is zero.
 Equation (1) above is seen as follows. First, it is convenient to note that (1) holds if
-$\lim\limits_{j\rightarrow\infty}\frac{(3t)^{j}}{2^{k_{j}d}}=0.$
-By taking logarithms, we see that (1) holds if
-$\lim\limits_{j\rightarrow\infty}\big{(}k_{j}d-j\log_{2}(3t)\big{)}=\infty.$
-As $d>0$, we let $c$ be the constant $(1/d)\log_{2}(3t)$ and we see that (1) holds if
-(2) $\lim\limits_{j\to\infty}(k_{j}-cj)=\infty.$
+$\lim\limits_{\ell\to\infty}\frac{(3t)^{c_{\ell}}}{2^{\ell s}}=0.$
+Now, by taking logarithms, we see that (1) holds if
+$\lim\limits_{\ell\to\infty}\Big{(}\ell s-c_{\ell}\log_{2}(3t)\Big{)}=\infty.$
+Let $\alpha=\frac{\log_{2}(3t)}{s}$. As $s>0$, we see that (1) holds if
+(2) $\lim\limits_{\ell\to\infty}\Big{(}\ell-\alpha c_{\ell}\Big{)}=\infty.$
 Now using Lemma 1, we have
-$j=g_{n}(k_{j})\leq(2+\log_{2}k_{j})^{n},$
+$c_{\ell}\leq(2+\log_{2}\ell)^{n},$
 which implies
-$k_{j}\geq 2^{\sqrt[n]{j}-2}.$
-Hence, (2) follows, which justifies our claim. ∎
+$\ell\geq 2^{c_{\ell}^{r}-2}$
+where $r=1/n$. Hence, it follows that (2) holds. This justifies our claim. ∎
 Let $n\in\mathbb{N}$.
@@
 $A:=\bigcup_{n=1}^{\infty}A_{n}.$
-We deduce that the set $A$ is an uncountable subgroup of $\mathbb{R}$ of Hausdorff dimension zero. We aim to establish that $A$ is in fact a subring. To this end, we need the following result.
+We deduce that the set AA is an uncountable subgroup of $\mathbb{R}$ of Hausdorff dimension zero. We aim to establish that AA is in fact a subring. To this end, we need the following result.
 Lemma 2.
-Let $n\in\mathbb{N}$ and let $a_{n}(s)$ be the number of representations of an element $s\in\mathbb{N}$ as a sum $s=t_{1}+t_{2}+\cdots+t_{n}$ of $n$ elements $t_{1},t_{2},\ldots,t_{n}\in S$.
-Then there exists $c_{n}\in\mathbb{R}$ such that $a_{n}(s)\leq c_{n}$ for every $s\in\mathbb{N}$.
+Let $n\in\mathbb{N}$ and let $c_{n}(k)$ be the number of representations of an element $k\in\mathbb{N}$ as a sum $k=s_{1}+s_{2}+\cdots+s_{n}$ of $n$ elements of $S$.
+Then there exists $b_{n}\in\mathbb{N}$ such that $c_{n}(k)\leq b_{n}$ for every $k\in\mathbb{N}$.
 Proof.
-Clearly $a_{1}(s)\leq 1$ for every $s\in\mathbb{N}$ and hence, we may take $c_{1}:=1$.
-Now assume that $n\geq 2$ and that there exists $c_{n-1}\in\mathbb{R}$ such that $a_{n-1}(s)\leq c_{n-1}$ for every $s\in\mathbb{N}$.
-Let $s\in\mathbb{N}$ and suppose
-$s=t_{1}+\cdots+t_{n}$
-is a representation of $s$ as a sum of $n$ elements of $S$.
-Then for at least one $i\in\{1,\ldots,n\}$, we have $t_{i}\geq s/n$ and hence $t_{i}\neq 0$.
-So there exists $k\in\mathbb{N}_{0}$ such that $t_{i}=2^{k}$.
-As $s/n\leq t_{i}\leq s$, we observe that $\log_{2}s-\log_{2}n\leq k\leq\log_{2}s$.
-So if
-$B_{n,s}:=[\,\log_{2}s-\log_{2}n,\,\log_{2}s\,]\cap\mathbb{N}_{0},$
+Clearly $c_{1}(k)\leq 1$ for every $k\in\mathbb{N}$ and hence, we may take $b_{1}:=1$.
+Now assume that $n\geq 2$ and that there exists $b_{n-1}\in\mathbb{N}$ such that $c_{n-1}(k)\leq b_{n-1}$ for every $k\in\mathbb{N}$.
+Let $k\in\mathbb{N}$ and suppose
+$k=s_{1}+\cdots+s_{n}$
+is a representation of $k$ as a sum of $n$ elements of $S$.
+Then for at least one $i\in\{1,\ldots,n\}$ we have $s_{i}\geq k/n$ and hence $s_{i}\neq 0$.
+So there exists $m\in\mathbb{N}_{0}$ such that $s_{i}=2^{m}$.
+As $k/n\leq s_{i}\leq k$, we observe that $\log_{2}k-\log_{2}n\leq m\leq\log_{2}k$.
+Let
+$I_{n,k}:=[\,\log_{2}k-\log_{2}n,\,\log_{2}k\,]\cap\mathbb{N}_{0}.$
 Then we see that
-$a_{n}(s)\leq n\sum\limits_{k\in B_{n,s}}a_{n-1}(s-2^{k})\leq n(1+\log_{2}n)c_{n-1}.$
-So we are done by taking $c_{n}:=n(1+\log_{2}n)c_{n-1}$. ∎
+$c_{n}(k)\leq n\sum\limits_{m\in I_{n,k}}c_{n-1}(k-2^{m})\leq nb_{n-1}\emph{{\#}}I_{n,k}\leq nb_{n-1}(1+\log_{2}n).$
+So we are done by taking $b_{n}\in\mathbb{N}$ such that $b_{n}\geq nb_{n-1}(1+\log_{2}n)$. ∎
 Theorem 3.
-The set $A$ is a a subring of $\mathbb{R}$.
+The set AA is a a subring of $\mathbb{R}$.
 Proof.
-It suffices to show that $1\in A$ and that $A$ is closed under multiplication. We see that
-$1=\frac{1}{2^{0}}+\sum\limits_{k\in S\setminus\{0\}}\frac{0}{2^{k}}\in A_{1}.$
-To establish that $A$ is closed under multiplication, we proceed as follows. Let $x,y\in A$. There exists $m,n\in\mathbb{N}$ such that $x\in A_{m}$ and $y\in A_{n}$. Therefore,
-$x=\sum\limits_{k\in mS}\frac{x_{k}}{2^{k}}\quad\mbox{and}\quad y=\sum\limits_{\ell\in nS}\frac{y_{\ell}}{2^{\ell}}$
-where $(x_{k})_{k\in mS}$ and $(y_{\ell})_{\ell\in nS}$ are bounded sequences of integers. Now
+It suffices to show that $1\in A$ and that AA is closed under multiplication. We see that
+$1=\frac{1}{2^{0}}+\sum\limits_{k\in S_{>0}}\frac{0}{2^{k}}\in A_{1}.$
+To establish that AA is closed under multiplication, we proceed as follows.
+Let $x,y\in A$.
+There exist $m,n\in\mathbb{N}$ such that $x\in A_{m}$ and $y\in A_{n}$.
+Therefore, there exist $t,t^{\prime}\in\mathbb{N}$ and for every $k\in nS$ there exist $x_{k}\in[-t,t\,]_{\mathbb{Z}}$ and $y_{k}\in[-t^{\prime},t^{\prime}]_{\mathbb{Z}}$ such that
+$x=\sum\limits_{k\in mS}\frac{x_{k}}{2^{k}}\quad\mbox{and}\quad y=\sum\limits_{\ell\in nS}\frac{y_{\ell}}{2^{\ell}}.$
+Now
 $xy=\left(\sum\limits_{k\in mS}\frac{x_{k}}{2^{k}}\right)\left(\sum\limits_{\ell\in nS}\frac{y_{\ell}}{2^{\ell}}\right)=\sum\limits_{(k,\ell)\in mS\times nS}\frac{x_{k}y_{\ell}}{2^{k+\ell}}=\sum\limits_{r\in(m+n)S}\frac{z_{r}}{2^{r}}$
+where for every $r\in(m+n)S$ we let
+$z_{r}:=\sum\limits_{\begin{subarray}{c}(k,\ell)\in mS\times nS\vskip 3.0pt plus 1.0pt minus 1.0pt\\ k+\ell=r\end{subarray}}x_{k}y_{\ell}.$
+The number of representations of an element $r\in(m+n)S$ as a sum $r=k+\ell$ of elements $k\in mS$ and $\ell\in nS$ is bounded by the number of representations of $r$ which are of the form $r=s_{1}+s_{2}+\cdots+s_{m+n}$ where $s_{1},s_{2},\ldots,s_{m+n}\in S$.
+Hence, by Lemma 2 we deduce that there exists $b\in\mathbb{N}$ such that for every $r\in(m+n)S$ we have $|z_{r}|\leq btt^{\prime}$.
+It follows that $xy\in A_{m+n}$.
+This establishes that AA is closed under multiplication.
+∎
+We now proceed to show that the only rational numbers in AA are the dyadic rationals. In particular, it will follow that the ring AA is not a field.
+Lemma 3.
+Let $b,n\in\mathbb{N}$. Then there exist two consecutive elements $k,k^{\prime}\in nS$ such that $k^{\prime}-k>b$.
+Proof.
+For every $\ell\in\mathbb{N}$ we let $c_{\ell}:=\emph{{\#}}\,(nS)_{\leq\ell}$. If the conclusion is false, then $c_{\ell}\geq\ell/b$ for every $\ell\in\mathbb{N}$. This contradicts Lemma 1. ∎
+Proposition 1.
+Let $x\in A$ and $\ell\in\mathbb{N}$. Then there exists a binary expansion of $x$ which contains either a string of zeroes or a string of ones having length $\ell$.
+Proof.
+Let $x\in A_{n,t}$. Then for every $k\in nS$ there exists $x_{k}\in[-t,t\,]_{\mathbb{Z}}$ such that
+$x=\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}.$
+Let $m\in nS$. We observe that
+$2^{m}x=q_{m}+r_{m}$
 where
-$z_{r}:=\sum\limits_{\begin{subarray}{c}(k,\ell)\in mS\times nS\\ k+\ell=r\end{subarray}}x_{k}y_{\ell}.$
-The number of representations of an element $r\in(m+n)S$ as a sum $r=k+\ell$ of elements $k\in mS$ and $\ell\in nS$ is bounded by the number of representations of $r$ which are of the form $r=t_{1}+t_{2}+\cdots+t_{m+n}$ where $t_{1},t_{2},\ldots,t_{m+n}\in S$.
-Hence, we deduce that the sequence $(z_{r})_{r\in(m+n)S}$ is bounded as a consequence of Lemma 2 and the boundedness of the sequences $(x_{k})_{k\in mS}$ and $(y_{\ell})_{\ell\in nS}$.
-It follows that $xy\in A_{m+n}$.
-This establishes that $A$ is closed under multiplication.
-∎
-We now proceed to show that the only rational numbers in $A$ are the dyadic rationals. In particular, it will follow that the ring $A$ is not a field.
-Lemma 3.
-Let $r,n\in\mathbb{N}$. Write $nS=\{k_{m}:m\in\mathbb{N}\}$ where $(k_{m})_{m\in\mathbb{N}}$ is an increasing sequence. Then there exists $m\in\mathbb{N}$ such that $k_{m+1}-k_{m}>r$.
-Proof.
-For every $z\in\mathbb{N}$, let $g_{n}(z):=\emph{{\#}}\{k\in nS:k\leq z\}$. If the conclusion is false, then $g_{n}(z)\geq z/r$ for every $z\in\mathbb{N}$. This contradicts Lemma 1. ∎
-Proposition 1.
-Let $x\in A$ and $l\in\mathbb{N}$. Then there exists a binary expansion of $x$ which contains either a string of zeroes or a string of ones having length $l$.
-Proof.
-Let $x\in A_{n,t}$. Then there exists a sequence of integers $(x_{k})_{k\in nS}$ with $|x_{k}|\leq t$ for all $k\in nS$ such that
-$x=\sum\limits_{k\in nS}\frac{x_{k}}{2^{k}}=\sum\limits_{i=1}^{\infty}\frac{x_{k_{i}}}{2^{k_{i}}}.$
-Let $m\in\mathbb{N}$. We observe that
-$2^{k_{m}}x=q_{m}+r_{m}$
-where
-$q_{m}:=\sum\limits_{i=0}^{m}2^{k_{m}-k_{i}}x_{k_{i}}\mbox{ and }r_{m}:=\sum\limits_{i=m+1}^{\infty}\frac{x_{k_{i}}}{2^{k_{i}-k_{m}}}.$
+$q_{m}:=\sum\limits_{k\in(nS)_{\leq m}}2^{m-k}x_{k}\mbox{ and }r_{m}:=\sum\limits_{k\in(nS)_{>m}}\frac{x_{k}}{2^{k-m}}.$
 Then $q_{m}\in\mathbb{Z}$ and we have
-$\left|r_{m}\right|\leq~{}t\sum\limits_{j=k_{m+1}-k_{m}}^{\infty}\frac{1}{2^{j}}=\frac{2t}{2^{k_{m+1}-k_{m}}}.$
-By Lemma 3, there exists $m\in\mathbb{N}$ such that $2^{k_{m+1}-k_{m}}>2^{l+1}t$ and so $|r_{m}|<1/2^{l}$. Hence,
-$\text{frac}\left(2^{k_{m}}x\right)=\begin{cases}r_{m},&\mbox{ if }r_{m}\in\left[\,0,\frac{1}{2^{l}}\right);\\ 1+r_{m},&\mbox{ if }r_{m}\in\left(-\frac{1}{2^{l}},0\right);\end{cases}$
+$|r_{m}|\leq~t\sum\limits_{j=m^{\prime}-m}^{\infty}\frac{1}{2^{j}}=\frac{2t}{2^{m^{\prime}-m}}$
+where $m,m^{\prime}\in nS$ are consecutive elements.
+By Lemma 3, there exist consecutive elements $k,k^{\prime}\in nS$ such that $2^{k^{\prime}-k}>2^{\ell+1}t$ and so we get $|r_{k}|<\frac{1}{2^{\ell}}$. Hence,
+$\text{frac}\left(2^{k}x\right)=\begin{cases}r_{k},&\mbox{ if }r_{k}\in\left[\,0,\frac{1}{2^{\ell}}\right);\\ 1+r_{k},&\mbox{ if }r_{k}\in\left(-\frac{1}{2^{\ell}},0\right);\end{cases}$
 and so
-$\text{frac}(2^{k_{m}}x)\in\left[0,\tfrac{1}{2^{l}}\right)\cup\left(1-\tfrac{1}{2^{l}},1\right),$
-where frac$(z):=z-[z]$ is the fractional part of $z\in\mathbb{R}$.
-It follows that the digits after the dot in a binary expansion of $x$ from the $(k_{m}+1)^{\text{th}}$ position to the $(k_{m}+l)^{\text{th}}$ position are all equal.
-∎
+$\text{frac}(2^{k}x)\in\left[0,\tfrac{1}{2^{\ell}}\right)\cup\left(1-\tfrac{1}{2^{\ell}},1\right),$
+[…]
 Corollary 1.
@@
 Proof.
-From Proposition 1, we see that the ring $A$ only contains those rational numbers which have a finite binary expansion. The result follows. ∎
+From Proposition 1, we see that the ring AA only contains those rational numbers which have a finite binary expansion. The result follows. ∎
 Remark 1.
 The above construction of an uncountable subring of $\mathbb{R}$ of Hausdorff dimension zero goes through if the initial set $S=\{0,1,2,2^{2},2^{3},\ldots\}$ is replaced by a set $T\cup\{0\}$ where $T$ is a subset of $\mathbb{N}$ which has the following property: There exists $b\in\mathbb{R}$ such that for every $n\in\mathbb{N}$, the cardinality of the set $\{\log t:t\in T\}\cap[n,n+1)$ is less than $b$.
-Acknowledgment: We thank Prof.
-Shrikrishna G.
-Dani for sharing with us a preprint of [1] in which we see a different construction of an uncountable subgroup of $\mathbb{R}$ having Hausdorff dimension zero.
+3. Concluding remarks
+In 1966, Erdős and Volkmann [2] showed that for every $d\in[0,1]$ there exists a subgroup of $\mathbb{R}$ which is Borel measurable and has Hausdorff dimension $d$.
+(Example 12.4 of [4]).
+In an unpublished work (see [3], pg.
+212), using the Continuum Hypothesis, R.
+O.
+Davies showed that for every $d\in[0,1]$ there exists a subring of $\mathbb{R}$ which has Hausdorff dimension $d$ and which is not a Borel set.
+In [1], it is shown that every proper subring of $\mathbb{R}$ which is a Borel set, has Hausdorff dimension zero.
+It has come to our attention that our construction has some similarities to that of [2], but our series expansion of real numbers is inspired by the binary expansion.
+Erdős and Volkmann [2] noted that the question of the Hausdorff dimension of subrings and subfields remained open.
+We feel that it is of interest to see how Lemma 2 is crucially used to show that we get a subring of $\mathbb{R}$ by using numbers which are represented by special series expansions.
+Also, we are able to see relatively easily that the set AA has Hausdorff dimension zero.
+We have been informed of the following construction.
+If AA is a compact Cantor set such that for every $n\in\mathbb{N}$ the set $A^{n}$ has Hausdorff dimension zero, then it can be shown that the subring of $\mathbb{R}$ which is generated by AA has Hausdorff dimension zero.
+However, we are unable to find a reference for the construction of such a set.
+Our construction is very explicit using series expansions and we use neither the Continuum Hypothesis nor the Axiom of choice.